OpenStudy (anonymous):

completing the sqaure of 0.5g^2+8g=-7 show work plez?!?

OpenStudy (anonymous):

multiply everything by 2 first of all g^2+16g=-14 take half of 16, which is 8, and square it. then add to each side g^2+16g+64=-14+64 (g+8)^2=50 if you want to solve, take square root of each side: \[g+8= \pm \sqrt{50}\] \[g+8= \pm 5\sqrt{2}\] \[g=-8 \pm 5\sqrt{2}\]

OpenStudy (campbell_st):

double each term g^2 + 16g = -14 for a quadratic ax^2 + bx + c = 0 find (b/2)^2 (16/2)^2 = 64 you need the constant to equal 64 g^2 + 16g + 14 + 50 = 0 (g + 8)^2 = 0

OpenStudy (campbell_st):

oops (g +8)^2 = 50